The carrying value of $60,000 face value bonds at the end of the third year is $61,116. The stated rate is 6%; the effective rate is 5%. On September 1, Year 4, the bonds are purchased on the market for $60,200 plus accrued interest. What is the gain or loss on retirement? Interest is paid annually.
Answer #1

Purchase date = sep 1st of year 4

Carrying value = $61,116 at the end of the third year .

since the compounding effect is considered we have taken the effective rate of interest 5% for calculation.

Accrued interest for 8 months i.e. = $2037 (61116 * 5% *8/12) ( 1st jan to 31st aug = 8 months)

Total value carrying value upon interest accural is end of 3 years and 8 months = $63153 i.e. $61,116 + $2037

Purchase value =Year 4 bonds acquired at $60200 + accrued interest , i.e. $60,200 + accrued interest $2,037 totals to $62,237

Termination value at the end of year 4 = 61,116 + 61,116*5% = 64,172

End of 4 years is considered as termination date

Gain would be = termination value at the end of year 4 less purchase value

=  $64,172 - $62,237

=$ 1935